The perimeter of a rectangle is 30 ft and its area is 54 ft^2. Find the dimensions.
Answer by sachi(548)

let the length of the rectangle be l
and breadth be b
ATQ 2(l+b)=30
so l+b=15...........eq^n 1
lb=54.............eq^n 2
so l-b=sqrt{(l+b)^2-4lb}=sqrt{15^2-4*54}=sqrt(225-216)=sqrt 9=3........eq^n 3
adding eq^n 1 & 2
2l=18
l=18/2=9
b=15-9=6 ans